OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is parallel to the line y = 3x +2 and passes through the point (5, 2). A. y = 3x – 5 B. y = 3x – 13 C. y = 3x – 1 D. y = 3x + 2

OpenStudy (anonymous):

It's B

OpenStudy (ash2326):

Given line y= 3x+2 standard equation of a line y=mx+c, m is slope, c is intercept for given line m=3 line parallel to it will also have a slope of 3 so let its equation be y=3x+c c? the line passes through (5,2) 2=3*5+c 2=15+c c=-13 y=3x-13 option B